# GraphTeam: Facilitating Large Language Model-based Graph Reasoning via Multi-Agent Collaboration

### v4

@@ 1 Introduction
Social networks Garton et al. (1997), urban computing Zheng et al. (2014), and various other fields make extensive use of graph data and related analytical methods. Given the promising problem-solving abilities demonstrated by large language models (LLMs) Brown et al. (2020); Anil et al. (2024); Anthropic (2024) across different domains, researchers are now exploring the feasibility of using LLMs to reason over graphs Tang et al. (2024); Sanford et al. (2024); Ouyang et al. (2024); Zhang et al. (2024a).
@@
Existing LLM-based graph reasoning methods mainly fall into two categories. The first category Chai et al. (2023); Liu et al. (2023); Qin et al. (2023); Liu et al. (2024) usually combines LLMs with graph neural networks (GNNs) for modeling. A typical paradigm is to first encode the input graph with a GNN, and then feed the encodings to an LLM as language token embeddings. However, this line of work is specialized for graph machine learning tasks, especially node classification, and thus has limited transferability. The second category Fatemi et al. (2024); Chen et al. (2024c) usually flattens graphs into textual descriptions, and purely relies on LLMs for analyzing Liu and Wu (2025); Huang et al. (2023); Zhao et al. (2024); Shi et al. (2023); Chen et al. (2024c). Strategies like in-context learning Brown et al. (2020) and chain-of-thought reasoning Wei et al. (2022) are widely used in these methods to improve performance. However, these methods analyze graphs by step-by-step reasoning on raw inputs without utilizing any external knowledge or tools, severely limiting their performance.
@@
Fortunately, LLMs have been shown to be capable of using external knowledge or tools for solving problems, known as a key feature of LLM-based agents Xi et al. (2023); Wang et al. (2024b). Moreover, inspired by the division of labor and cooperation mechanisms in human society, researchers propose LLM-based multi-agent systems Qian et al. (2024); Hong et al. (2024); Wu et al. (2024); Chen et al. (2024b), where agents with different specialities can collaborate with each other to address complex problems. But previous multi-agent frameworks are designed for general purposes, and have few knowledge relevant to graph reasoning.
@@
Inspired by human problem-solving processes Robson and Pólya (1945); Eisenberg and Berkowitz (1990); Bybee et al. (2006); Ramnarine-Rieks (2011), we present GraphTeam, a multi-agent system with three core modules specialized for graph reasoning: (1) Input-output normalization module. In human problem solving, clearly understanding and formalizing the problem is the first step Robson and Pólya (1945); Bybee et al. (2006), and answering the question in required formats is also important. Therefore, we introduce two different LLM-based agents: the question agent extracts and refines key arguments from the original question to help the system better understand the problem; the answer agent reformats the results to meet the specified output requirement. (2) External knowledge retrieval module. In tackling complex problems, humans often seek help from external sources Eisenberg and Berkowitz (1990) or draw analogies to previously solved ones Bybee et al. (2006). Therefore, we first build a knowledge base that includes relevant documentation and experience information. Then we introduce a search agent to retrieve the most relevant entries from this knowledge base to assist the downstream solving process. (3) Problem-solving module. Given the retrieved information, we introduce a coding agent that uses established algorithms via programming to generate solutions. If the coding agent fails, a reasoning agent directly infers the solution without relying on programming. The overall pipeline of GraphTeam is shown in Figure 1. We evaluate GraphTeam through extensive experiments across six graph reasoning benchmarks. These benchmarks cover a variety of graph tasks, ranging from basic graph theory to the deployment of GNNs. The experimental results and ablation studies highlight the effectiveness of GraphTeam as well as the necessity of each module. In summary, our contributions are three-fold:
@@ 2.1 LLM-based Graph Reasoning
To evaluate the ability of LLMs in solving graph reasoning problems, researchers have developed various benchmarks Guo et al. (2023); Fatemi et al. (2024); Das et al. (2024); Chen et al. (2024a); Li et al. (2024b). Among them, NLGraph Wang et al. (2024a) is one of the first benchmark for assessing LLMs’ graph analysis capabilities, focusing on basic graph theory reasoning and simple GNN computations. Graphwiz Chen et al. (2024a) evaluates LLMs on graph problems of varying complexity, including NP-complete ones. GraphInstruct Luo et al. (2024) includes 21 graph analysis tasks covering node-level, node-pair-level, and graph-level problems. Talk like a Graph Fatemi et al. (2024) measures basic graph understanding abilities, while LLM4DyG Zhang et al. (2024b) focuses on dynamic graph tasks related to temporal and spatial information. In our experiments, we employ the above five benchmarks, and also construct an additional one to evaluate the ability of deploying GNNs with AutoGL Guan et al. (2021).
There is a rising trend of research to utilize LLMs for various graph reasoning tasks Li et al. (2024c); Jin et al. (2024). Here we focus on the most relevant methods that employ LLMs as predictors Chai et al. (2023); Liu et al. (2023); Qin et al. (2023); Liu et al. (2024). Related work can be roughly divided into two categories. The first category typically integrates LLMs with graph neural network (GNNs), and is specialized for graph machine learning tasks such as node classification Tang et al. (2024); Zhang et al. (2024a). For example, GraphGPT Tang et al. (2024) employs instruction tuning to fine-tune a GNN-enhanced LLM for node classification. GraphTranslator Zhang et al. (2024a) encodes nodes in a graph by a GNN, and then project node embeddings into the embedding space of LLM tokens for zero-shot classification. However, these methods are designed and optimized for a specific task, and cannot generalize well to the diverse reasoning tasks in the above benchmarks. The second category usually flatten graphs into textual descriptions of adjacency lists, and then rely on LLM itself to analyze graphs with strategies like in-context learning Brown et al. (2020), chain-of-thought reasoning Wei et al. (2022) for inference Liu and Wu (2025); Huang et al. (2023); Zhao et al. (2024); Shi et al. (2023); Chen et al. (2024c). For example, GUNDAM Ouyang et al. (2024) collects natural language-based question-answer pairs of graph analysis problems to fine-tune LLMs, and performs step-by-step reasoning for inference. However, recent studies have shown that the reasoning depths of current LLMs are still shallow Parmar et al. (2024); Li et al. (2024a), and will probably fail for problems that require complex reasoning. In contrast, we encourage LLMs to leverage external knowledge and tools, instead of performing step-by-step reasoning on raw inputs.
@@ 2.2 Multi-Agent Collaboration of LLMs
Autonomous agents aim to solve problems with self-directed planning and actions. Recently, researchers introduce the concept of LLM-based agent that is capable of understanding natural language instruction, interacting with humans, perceiving external environments, and performing various actions Bang et al. (2023); Li et al. (2023); Xi et al. (2023); Wang et al. (2024b). Moreover, inspired by the collaboration of human in solving problems, researchers introduce LLM-based multi-agent systems, where agents collaborate with a well-defined division of labor. For example, ChatDev Qian et al. (2024) is a virtual software company where LLM-based agents are responsible for different software development tasks (such as user interface design or software testing) for collaboration. MetaGPT Hong et al. (2024) is a meta-programming framework that simulates human workflows. With proper developing pipeline, MetaGPT can generate effective programming solutions. AgentVerse Chen et al. (2024b) is a multi-agent framework inspired by human collaboration dynamics. It involves four iterative stages of expert recruitment, decision discussion, action and evaluation. AutoGen Wu et al. (2024) is another multi-agent framework for general purposes. Through tool invocation and code generation, AutoGen can accomplish a wide variety of tasks. MACRec Wang et al. (2024c) is a novel framework designed to enhance recommendation systems through multi-agent collaboration. However, these multi-agent frameworks are not specialized for graph reasoning, e.g., they have few knowledge of relevant Python libraries and no experience base of similar problems.
@@ 3.2 Question Agent
Abstraction is a preliminary step for human beings to solve problems Robson and Pólya (1945). Inspired by this, we introduce a question agent to extract and refine four key arguments from the original problem descriptions. Formally, given a graph reasoning problem with its original description denoted as $q$, we employ an LLM denoted as $\operatorname{Question}(q)$ to generate refined question $q_{r}$, graph type $q_{t}$, input graph $q_{g}$ and output format $q_{f}$:
@@ 3.3 Search Agent
During the solving process of a hard problem, it is common for human beings to seek help from external sources Eisenberg and Berkowitz (1990) or make analogy to a previously solved one Bybee et al. (2006). Therefore, we first build a knowledge base with documentation and experience information, and then ask the search agent to retrieve relevant entries from the knowledge base for each question.
@@ 3.3.2 Knowledge Retrieval
For effective information retrieval, we build vector databases with LlamaIndex Liu (2022) for documentation and experiences, respectively. Then the search agent will take the refined question $q_{r}$ and graph type $q_{t}$ as query, and find the most relevant experience and API information from the database. Note that the experiences of similar problems contains detailed solutions, and are usually more helpful for downstream agents than documentations. Therefore, we first use similarity matching via LlamaIndex to score the experiences. If the largest similarity score exceeds a predefined threshold $\delta$, the corresponding experience will be directly returned by the search agent without querying related documentation information. Otherwise, the search agent will return the most related entries in documentation base. Finally, either the experience or the documentation information returned by the search agent will be passed to the downstream agents for retrieval-augmented generation (RAG). For simplicity, we denote the operation of the search agent as
@@ 3.4 Coding and Reasoning Agents
Given the outputs of question and search agents, the coding agent will generate and execute Python codes to solve the problems. In practice, the generated codes may encounter compilation or running errors. Following the trial-and-error strategy Robson and Pólya (1945); Ramnarine-Rieks (2011) for problem solving, we introduce a retry mechanism that attempts to fix the codes with previous codes and error messages. More formally, the operation of coding agent in the $n$-th trial can be written as:
@@
where $\text{result}_{q}^{n},\text{code}_{q}^{n},\text{error}_{q}^{n}$ are respectively the execution result, generated code, and error message of the $n$-th trial. The last two parameters are the codes and error messages from previous $n-1$ trials. If the codes runs normally, we define $\text{error}_{q}^{n}$ as $\tt None$.
@@ 3.6 Summary
We formalize the inference pipeline of GraphTeam in Alg. 1. Most agent functions (including Question, Coding, Reasoning and Answer) in Alg. 1 are all implemented by an LLM with different prompt templates. The Search function is implemented based on LlamaIndex Liu (2022). We will provide more implementation details in Appendix A.
@@
1:  $(q_{r},q_{t},q_{g},q_{f})\leftarrow\operatorname{Question}(q)$
2:  $\text{knowledge}_{q}\leftarrow\operatorname{Search}(q_{r},q_{t},\mathcal{K},\delta)$
@@ 4.1.1 Datasets
To evaluate the effectiveness of our system, we consider six benchmarks for LLM-based graph reasoning: Talk like a Graph Fatemi et al. (2024) primarily studies some fundamental problems in graph reasoning and explores the cognitive abilities of LLMs regarding graphs. LLM4DyG Zhang et al. (2024b) focuses on basic problems related to dynamic graphs, while GraphWiz Chen et al. (2024a), NLGraph Wang et al. (2024a), and GraphInstruct Luo et al. (2024) are used to evaluate the understanding and mastery of algorithms in fundamental graph theory. GNN-AutoGL is a benchmark we constructed in this paper to study the ability of LLMs in deploying GNNs with AutoGL Guan et al. (2021). Dataset statistics and details can be found at Appendix B.1 and  B.2.
@@ 4.1.2 Metrics
For benchmarks with definite answers Fatemi et al. (2024); Wang et al. (2024a); Zhang et al. (2024b); Chen et al. (2024a); Luo et al. (2024), we use accuracy as the evaluation metric. For GNN-AutoGL, since the task requires a complex pipeline (including identifying proper APIs, filling correct hyper-parameters, successfully training and evaluating the corresponding GNN model on a specified dataset, we evaluate whether the generated codes can be executed correctly and whether the key hyper-parameters (e.g., dataset name, model name, hidden dimension, dropout) are correct as the evaluation criteria. More details are shown in Appendix B.3.
@@ 4.1.3 Settings
We choose GPT-4o-mini Hurst et al. (2024) and Qwen2.5-Coder-7B Hui et al. (2024) as the base LLM for the agents of GraphTeam. For Python libraries, we use NetworkX version 3.3 Hagberg et al. (2008) and AutoGL version 0.4.0 Guan et al. (2021), together with Python version 3.10.14. The hyper-parameters are set as follows: the number of candidate experiences $N_{exp}=10$, similarity matching threshold for using experience $\delta=0.85$. The maximum number of retries for the coding agent is $3$, and the number of self-checking iterations for the answer agent is $3$.
@@ 4.1.4 Baselines
As the SOTA methods for each benchmark are quite diverse and some of which are designed for a specific benchmark, we directly compare with the previously reported SOTA performance for each benchmark: for NLGraph, we select GUNDAM Ouyang et al. (2024) as the SOTA; for Talk like a Graph, we select the fine-tuned transformers Sanford et al. (2024) as the SOTA; for LLM4DyG, GraphWiz, and GraphInstruct Zhang et al. (2024b); Chen et al. (2024a); Luo et al. (2024), we select the best model in their benchmark papers. In addition, we adopt AutoGPT Significant Gravitas (2023) and Open Interpreter Open Interpreter (2023) as single-agent baselines, while employing AutoGen Wu et al. (2024) and ChatDev Qian et al. (2024) as multi-agent baselines for comparison.
@@ 4.2 Main Results
Base LLM       Methods NLGraph Talk like a Graph GraphInstruct GraphWiz LLM4DyG GNN-AutoGL Average Rank
@@ 4.3 Ablation Study
Category Methods NLGraph Talk like a Graph GraphInstruct GraphWiz LLM4DyG GNN-AutoGL
@@
Input-Output Normalization: For most benchmark problems, the question agent significantly contributes (25.3%) by breaking the problem into individual components on average. For a few benchmark such as GraphInstruct Luo et al. (2024), where the structure of the question is already clear and straightforward, the impact of the question agent is less noticeable (less than 2.2%). The answer agent’s correction of the format is also an essential part of correctly answering the questions (0.6-2.8%).

### v5

@@ 1 Introduction
Social networks [13], urban computing [51], and various other fields make extensive use of graph data and related analytical methods. Given the promising problem-solving abilities demonstrated by large language models (LLMs) [4, 1, 2] across different domains, researchers are now exploring the feasibility of using LLMs to reason over graphs [41, 38, 32, 48].
@@
Existing LLM-based graph reasoning methods mainly fall into two categories. The first category [6, 29, 35, 28] usually combines LLMs with graph neural networks (GNNs) for modeling. A typical paradigm is to first encode the input graph with a GNN, and then feed the encodings to an LLM as language token embeddings. However, this line of work is specialized for graph machine learning tasks, especially node classification, and thus has limited transferability. The second category [12, 9] usually flattens graphs into textual descriptions, and purely relies on LLMs for analyzing [26, 18, 50, 39, 9]. Strategies like in-context learning [4] and chain-of-thought reasoning [45] are widely used in these methods to improve performance. However, these methods analyze graphs by step-by-step reasoning on raw inputs without utilizing any external knowledge or tools, severely limiting their performance.
@@
Fortunately, LLMs have been shown to be capable of using external knowledge or tools for solving problems, known as a key feature of LLM-based agents [47, 43]. Moreover, inspired by the division of labor and cooperation mechanisms in human society, researchers propose LLM-based multi-agent systems [34, 17, 46, 8], where agents with different specialities can collaborate with each other to address complex problems. But previous multi-agent frameworks are designed for general purposes, and have few knowledge relevant to graph reasoning.
@@
Inspired by human problem-solving processes [37, 11, 5, 36], we present GraphTeam, a multi-agent system with three core modules specialized for graph reasoning: (1) Input-output normalization module. In human problem solving, clearly understanding and formalizing the problem is the first step [37, 5], and answering the question in required formats is also important. Therefore, we introduce two different LLM-based agents: the question agent extracts and refines key arguments from the original question to help the system better understand the problem; the answer agent reformats the results to meet the specified output requirement. (2) External knowledge retrieval module. In tackling complex problems, humans often seek help from external sources [11] or draw analogies to previously solved ones [5]. Therefore, we first build a knowledge base that includes relevant documentation and experience information. Then we introduce a search agent to retrieve the most relevant entries from this knowledge base to assist the downstream solving process. (3) Problem-solving module. Given the retrieved information, we introduce a coding agent that uses established algorithms via programming to generate solutions. If the coding agent fails, a reasoning agent directly infers the solution without relying on programming. The overall pipeline of GraphTeam is shown in Figure LABEL:fig:simplified_modules_and_overall_performance. We evaluate GraphTeam through extensive experiments across six graph reasoning benchmarks. These benchmarks cover a variety of graph tasks, ranging from basic graph theory to the deployment of GNNs. The experimental results and ablation studies highlight the effectiveness of GraphTeam as well as the necessity of each module. In summary, our contributions are three-fold:
@@ 2.1 LLM-based Graph Reasoning
To evaluate the ability of LLMs in solving graph reasoning problems, researchers have developed various benchmarks [15, 12, 10, 7, 24]. Among them, NLGraph [42] is one of the first benchmark for assessing LLMs’ graph analysis capabilities, focusing on basic graph theory reasoning and simple GNN computations. Graphwiz [7] evaluates LLMs on graph problems of varying complexity, including NP-complete ones. GraphInstruct [30] includes 21 graph analysis tasks covering node-level, node-pair-level, and graph-level problems. Talk like a Graph [12] measures basic graph understanding abilities, while LLM4DyG [49] focuses on dynamic graph tasks related to temporal and spatial information. In our experiments, we employ the above five benchmarks, and also construct an additional one to evaluate the ability of deploying GNNs with AutoGL [14].
There is a rising trend of research to utilize LLMs for various graph reasoning tasks [25, 21]. Here we focus on the most relevant methods that employ LLMs as predictors [6, 29, 35, 28]. Related work can be roughly divided into two categories. The first category typically integrates LLMs with graph neural network (GNNs), and is specialized for graph machine learning tasks such as node classification [41, 48]. For example, GraphGPT [41] employs instruction tuning to fine-tune a GNN-enhanced LLM for node classification. GraphTranslator [48] encodes nodes in a graph by a GNN, and then project node embeddings into the embedding space of LLM tokens for zero-shot classification. However, these methods are designed and optimized for a specific task, and cannot generalize well to the diverse reasoning tasks in the above benchmarks. The second category usually flatten graphs into textual descriptions of adjacency lists, and then rely on LLM itself to analyze graphs with strategies like in-context learning [4], chain-of-thought reasoning [45] for inference [26, 18, 50, 39, 9]. For example, GUNDAM [32] collects natural language-based question-answer pairs of graph analysis problems to fine-tune LLMs, and performs step-by-step reasoning for inference. However, recent studies have shown that the reasoning depths of current LLMs are still shallow [33, 23], and will probably fail for problems that require complex reasoning. In contrast, we encourage LLMs to leverage external knowledge and tools, instead of performing step-by-step reasoning on raw inputs.
@@ 2.2 Multi-Agent Collaboration of LLMs
Autonomous agents aim to solve problems with self-directed planning and actions. Recently, researchers introduce the concept of LLM-based agent that is capable of understanding natural language instruction, interacting with humans, perceiving external environments, and performing various actions [3, 22, 47, 43]. Moreover, inspired by the collaboration of human in solving problems, researchers introduce LLM-based multi-agent systems, where agents collaborate with a well-defined division of labor. For example, ChatDev [34] is a virtual software company where LLM-based agents are responsible for different software development tasks (such as user interface design or software testing) for collaboration. MetaGPT [17] is a meta-programming framework that simulates human workflows. With proper developing pipeline, MetaGPT can generate effective programming solutions. AgentVerse [8] is a multi-agent framework inspired by human collaboration dynamics. It involves four iterative stages of expert recruitment, decision discussion, action and evaluation. AutoGen [46] is another multi-agent framework for general purposes. Through tool invocation and code generation, AutoGen can accomplish a wide variety of tasks. MACRec [44] is a novel framework designed to enhance recommendation systems through multi-agent collaboration. However, these multi-agent frameworks are not specialized for graph reasoning, e.g., they have few knowledge of relevant Python libraries and no experience base of similar problems.
@@ 3.2 Question Agent
Abstraction is a preliminary step for human beings to solve problems [37]. Inspired by this, we introduce a question agent to extract and refine four key arguments from the original problem descriptions. Formally, given a graph reasoning problem with its original description denoted as $q$, we employ an LLM denoted as $\operatorname{Question}(q)$ to generate refined question $q_{r}$, graph type $q_{t}$, input graph $q_{g}$ and output format $q_{f}$:
@@ 3.3 Search Agent
During the solving process of a hard problem, it is common for human beings to seek help from external sources [11] or make analogy to a previously solved one [5]. Therefore, we first build a knowledge base with documentation and experience information, and then ask the search agent to retrieve relevant entries from the knowledge base for each question.
@@ 3.3.2 Knowledge Retrieval
For effective information retrieval, we build vector databases with LlamaIndex [27] for documentation and experiences, respectively. Then the search agent will take the refined question $q_{r}$ and graph type $q_{t}$ as query, and find the most relevant experience and API information from the database. Note that the experiences of similar problems contains detailed solutions, and are usually more helpful for downstream agents than documentations. Therefore, we first use similarity matching via LlamaIndex to score the experiences. If the largest similarity score exceeds a predefined threshold $\delta$, the corresponding experience will be directly returned by the search agent without querying related documentation information. Otherwise, the search agent will return the most related entries in documentation base. Finally, either the experience or the documentation information returned by the search agent will be passed to the downstream agents for retrieval-augmented generation (RAG). For simplicity, we denote the operation of the search agent as
@@ 3.4 Coding and Reasoning Agents
Given the outputs of question and search agents, the coding agent will generate and execute Python codes to solve the problems. In practice, the generated codes may encounter compilation or running errors. Following the trial-and-error strategy [37, 36] for problem solving, we introduce a retry mechanism that attempts to fix the codes with previous codes and error messages. More formally, the operation of coding agent in the $n$-th trial can be written as:
@@
where $\text{result}_{q}^{n},\text{code}_{q}^{n},\text{error}_{q}^{n}$ are respectively the execution result, generated code, and error message of the $n$-th trial. The last two parameters are the codes and error messages from previous n1n-1 trials. If the codes runs normally, we define $\text{error}_{q}^{n}$ as $\tt None$.
@@ 3.6 Summary
We formalize the inference pipeline of GraphTeam in Alg. 1. Most agent functions (including Question, Coding, Reasoning and Answer) in Alg. 1 are all implemented by an LLM with different prompt templates. The Search function is implemented based on LlamaIndex [27]. We will provide more implementation details in Appendix A.
@@
[…]
@@ 4.1.1 Datasets
To evaluate the effectiveness of our system, we consider six benchmarks for LLM-based graph reasoning: Talk like a Graph [12] primarily studies some fundamental problems in graph reasoning and explores the cognitive abilities of LLMs regarding graphs. LLM4DyG [49] focuses on basic problems related to dynamic graphs, while GraphWiz [7], NLGraph [42], and GraphInstruct [30] are used to evaluate the understanding and mastery of algorithms in fundamental graph theory. GNN-AutoGL is a benchmark we constructed in this paper to study the ability of LLMs in deploying GNNs with AutoGL [14]. Dataset statistics and details can be found at Appendix B.1 and  B.2.
@@ 4.1.2 Metrics
For benchmarks with definite answers [12, 42, 49, 7, 30], we use accuracy as the evaluation metric. For GNN-AutoGL, since the task requires a complex pipeline (including identifying proper APIs, filling correct hyper-parameters, successfully training and evaluating the corresponding GNN model on a specified dataset, we evaluate whether the generated codes can be executed correctly and whether the key hyper-parameters (e.g., dataset name, model name, hidden dimension, dropout) are correct as the evaluation criteria. More details are shown in Appendix B.3.
@@ 4.1.3 Settings
We choose GPT-4o-mini [20] and Qwen2.5-Coder-7B [19] as the base LLM for the agents of GraphTeam. For Python libraries, we use NetworkX version 3.3 [16] and AutoGL version 0.4.0 [14], together with Python version 3.10.14. The hyper-parameters are set as follows: the number of candidate experiences $N_{exp}=10$, similarity matching threshold for using experience $\delta=0.85$. The maximum number of retries for the coding agent is $3$, and the number of self-checking iterations for the answer agent is $3$.
@@ 4.1.4 Baselines
As the SOTA methods for each benchmark are quite diverse and some of which are designed for a specific benchmark, we directly compare with the previously reported SOTA performance for each benchmark: for NLGraph, we select GUNDAM [32] as the SOTA; for Talk like a Graph, we select the fine-tuned transformers [38] as the SOTA; for LLM4DyG, GraphWiz, and GraphInstruct [49, 7, 30], we select the best model in their benchmark papers. In addition, we adopt AutoGPT [40] and Open Interpreter [31] as single-agent baselines, while employing AutoGen [46] and ChatDev [34] as multi-agent baselines for comparison.
@@ 4.2 Main Results
Base LLM     Methods NLGraph
Talk like a Graph
GraphInstruct GraphWiz LLM4DyG GNN-AutoGL Average Rank
@@ 4.3 Ablation Study
Category Methods NLGraph
Talk like a Graph
GraphInstruct GraphWiz LLM4DyG GNN-AutoGL
@@
Input-Output Normalization: For most benchmark problems, the question agent significantly contributes (25.3%) by breaking the problem into individual components on average. For a few benchmark such as GraphInstruct [30], where the structure of the question is already clear and straightforward, the impact of the question agent is less noticeable (less than 2.2%). The answer agent’s correction of the format is also an essential part of correctly answering the questions (0.6-2.8%).